# Prompting in Practice: Investigating Software Developers’ Use of Generative AI Tools

### v1

@@ Abstract.
The integration of generative artificial intelligence (GenAI) tools has fundamentally transformed software development. Although prompt engineering has emerged as a critical skill, existing research focuses primarily on individual techniques rather than software developers’ broader workflows. This study presents a systematic investigation of how software engineers integrate GenAI tools into their professional practice through a large-scale survey examining prompting strategies, conversation patterns, and reliability assessments across various software engineering tasks.
We surveyed 91 software engineers, including 72 active GenAI users, to understand AI usage patterns throughout the development process. Our ???? key findings show that while code generation is nearly universal, proficiency strongly correlates with using AI for more nuanced tasks such as debugging and code review, and that developers prefer iterative multi-turn conversations to single-shot prompting. Documentation tasks are perceived as most reliable, while complex code generation and debugging present sizable challenges. Our insights provide an empirical baseline of current developer practices, from simple code generation to deeper workflow integration, with actionable insights for future improvements.
@@ 1. Introduction
Generative Artificial Intelligence (GenAI) is rapidly transforming the way developers create and evolve software. With the help of GenAI tools such as GitHub Copilot (GitHub, [n. d.]), Windsurf (codeium, [n. d.]), and ChatGPT (OpenAI, [n. d.]), developers can now easily and quickly generate new code, review and modify existing codebases, identify and correct defects, and produce code documentation. Central to these capabilities is prompting: the practice of crafting natural language instructions to guide GenAI systems toward producing useful and accurate output. Prompting has emerged as a novel engineering practice, fundamentally different from traditional software development due to its reliance on large language models (LLMs), which are inherently probabilistic and non-deterministic  (Atil et al., 2025).
@@
Previous research has primarily focused on cataloging prompting techniques (White et al., 2023; Schulhoff et al., 2025; Sahoo et al., 2025), which offers limited insight into how developers actually use prompting in practice: how they craft prompts for different types of software engineering (SE) tasks, how they adapt to unreliable or suboptimal output, and how GenAI is reshaping standard development activities. Previous studies have examined narrow use cases (e.g., refactoring (AlOmar et al., 2024)), specific populations (e.g., students (Alpizar-Chacon and Keuning, 2025; Choudhuri et al., 2025a)), and broader socio-organizational adoption factors (Li et al., 2024; Russo, 2024), but there remains a lack of systematic investigation into real-world prompting practices in the full range of SE tasks.
@@
To address this gap, we conducted a comprehensive survey of 72 software developers, exploring how they use GenAI tools and engage in prompting across six common SE tasks: code generation, documentation, debugging, testing, refactoring, and review. Our study investigates various dimensions of GenAI usage, including the prompting and conversational strategies developers employ, how they assess and respond to the quality of GenAI outputs, and how prompting practices are integrated into their daily workflows. Together, these dimensions provide a holistic view of how prompting functions in modern software development and extend the current understanding of prompt engineering in practice.
@@
We find that while nearly all developers who use GenAI apply it to code generation, higher perceived proficiency strongly correlates with engagement in more types of tasks, such as code refactoring and testing. Developers rely primarily on iterative conversations, with daily usage critical to achieving significant productivity improvements. Our reliability assessment reveals that developers trust GenAI most for documentation and least for complex code generation/modification tasks. These findings have important implications for developers seeking to improve productivity through effective prompting strategies, for tool builders designing more reliable and intuitive GenAI, and for educators developing curricula around the evolving skill set needed for effective GenAI collaboration.
@@
This paper makes the following contributions: (1) a comprehensive characterization of how GenAI tools are used in various SE tasks; (2) empirical insights into developers’ prompting strategies, reliability perceptions, and common failure types; and (3) actionable implications for improving human-AI interaction in SE.
@@ 2. Study Methodology
The goal of our study is to investigate how developers use prompt engineering and GenAI tools for development tasks and what challenges they face during this use. The context consists of software developers with experience using GenAI tools for SE.
@@
    How do software developers integrate GenAI tools into their development workflows?
@@
    What prompting techniques and conversation strategies do developers use when interacting with GenAI?
@@
    How reliable do developers perceive GenAI to be for various SE tasks and what are common issues they encounter?
Below, we detail our methodology (Figure 1), which aligns with our prior studies (Stalnaker et al., 2024a, b; Wintersgill et al., 2024; Zappin et al., 2025) and was approved by our institution’s ethics review board, including the questionnaire, participant identification procedure, and data collection/distribution.
Figure 1. Study Methodology (image credits at (Otten et al., 2025))
@@ 2.1. Survey Design
We designed our survey questionnaire considering existing literature on prompt engineering and GenAI tool use (White et al., 2023; Schulhoff et al., 2025), following general guidelines (Groves et al., 2009) and SE-specific guidelines for survey design (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). Through the survey, our goal was to build an understanding of GenAI integration across six core SE tasks: code generation, refactoring, testing, debugging, documentation, and code review. Although other SE tasks exist, we opted to include only these six in an attempt to balance survey length and appropriate task coverage. We also implemented branching logic in the survey, so that participants were only shown questions relevant to their experiences. To ensure data quality, we also included an initial question to screen out respondents with no experience using prompt-based GenAI tools. We carefully formulated survey questions, ensuring that they were written clearly to avoid confusing or biasing respondents. We also conducted a pilot study with graduate students from our research lab who regularly use GenAI tools. Based on their feedback, we further refined our questionnaire, improving questions that were unclear or confusing.
The survey was organized into five sections. Section 1 collected demographic information. Sec. 2 asked questions related to the GenAI experience of the participants, the GenAI tools they regularly use, and their GenAI development workflows. Sec. 3 and 4 investigated the use and familiarity of participants with prompting techniques and conversation strategies. The final section explored the perceived reliability of GenAI tools and the issues that developers face. Questions were presented in a variety of formats, including multiple-selection, Likert scales, and open text fields. The full text and logic of our survey can be found in our replication package (Otten et al., 2025).
@@ 2.2. Participant Recruitment
Our recruitment strategy targeted developers with diverse backgrounds and experience levels. For survey distribution, we mainly relied on our professional networks and snowball sampling (e.g., participants sharing the survey within their networks) by directly contacting potential participants or posting participation calls in social media platforms (e.g., LinkedIn and X). To further broaden our sample and include participants at earlier career stages, we distributed the survey to relevant organizations at our university, including the Association for Computing Machinery (ACM) chapter. Participants were not compensated monetarily or otherwise.
@@ 2.3. Data Collection and Analysis
Survey responses were collected using Qualtrics (qua, nd) between 27 May and 7 July 2025. To analyze the data, we employed a mixed method approach, combining quantitative and qualitative techniques.
For multiple choice and Likert scale questions, we calculated data distributions and descriptive statistics to identify trends and patterns. Were appropriate, to facilitate comparison, we assigned ordinal Likert scale responses (e.g., ”Often” to ”Never”) to integer values (e.g., 1 to 4). This allowed us to compute mean scores to represent concepts such as the popularity of a prompting strategy or the frequency of an issue using GenAI tools.
The responses to open-ended text responses were analyzed using a qualitative coding approach (Spencer, 2009). Three annotators (i.e., authors) independently analyzed open-ended text responses and performed open-coding by assigning one or more codes to each response using a shared codebook (i.e., a spreadsheet). After the initial coding phase, the annotators met to collaboratively review, discuss and reconcile the coding decisions. Through this process, the annotators resolved any discrepancies, which included differing interpretations of potentially ambiguous responses, applying the same codes in different contexts, and using broader codes rather than more specific ones. Ultimately, this process resulted in a final set of codes, definitions, and labels for each response, which are found in our replication package (Otten et al., 2025). Since codes were developed iteratively through an inductive process and multiple codes could be assigned to each response, we did not base our analysis on inter-annotator agreements. We followed best open-coding practices (Spencer, 2009) and used annotator discussions to ensure the reliability of the results.
@@ 2.4. Participant Demographics
The 72 participants represent a diverse, global sample of software professionals, with reported experience ranging from one to 40 years (mean = 10.2, median = 10). They held a variety of roles, including researcher (64%), programmer (60%), technical lead (28%), and software architect (28%). Respondents could select multiple roles, and overlaps were common: for example, while 46 participants identified as researchers (the most frequently selected role) 31 of them also reported other roles, notably programmer, architect, technical lead, and/or tester. For comparative analysis, we grouped professional experience into four categories: 24 juniors (1–5 years), 22 intermediates (6–10 years), 19 seniors (11–20 years), and 7 veterans (21+ years).
@@ 3.1.1. GenAI Use for SE Tasks and Proficiency
Table 1 shows that the vast majority of respondents (91.7%) regularly use GenAI for code generation. Although this is somewhat expected, we found that debugging (47.2%) and code documentation (44.4%) emerged as next most common tasks. The remaining tasks, testing (38.9%, refactoring (31.9%), and code review (31.9%) were less frequently reported. Overall, 89% of the respondents reported using GenAI for four or fewer tasks, with two (30.6%) or three (23.6%) tasks being most common, suggesting a relatively focused use of GenAI.
When analyzing the 5-Likert point responses of self-reported proficiency in using GenAI, we found that the majority of participants (57 of 72) considered themselves proficient or very/maximally proficient (see Table 1). (Since a single veteran developer rated themselves as having “minimum proficiency”, we exclude this data point from future comparisons across proficiency levels.)
We also observed that proficiency correlates with the number of SE tasks for which participants use GenAI.(Spearman’s $\rho$ = 0.325). Among those who reported using GenAI for four or more SE tasks, most (18 of 22, 82%) identified as very or maximally proficient. By contrast, single-task users reported being somewhat proficient or proficient (7 of 11, 64%). These results suggest that task breadth and GenAI proficiency may mutually reinforce one another.
When examining proficiency by software development experience, we found that junior and intermediate developers considered themselves more proficient in using GenAI that seniors and veterans. Specifically, 22 of 24 juniors and 19 of 22 intermediates reported being proficient or very/maximally proficient, compared to only 12 of 19 seniors and 4 of 7 veterans.
Finding 1: A moderate correlation between perceived GenAI proficiency and GenAI use across six SE tasks suggests that that greater proficiency is associated with broader usage, and vice versa. Junior and intermediate developers tend to rate themselves as more proficient in using GenAI compared to senior and veteran developers.
@@ 3.1.2. GenAI Usage Frequency and Tool Adoption
Respondents who reported using GenAI more frequently tended to also report using GenAI for more SE task and experiencing productivity gains (Figure 2). Daily users reported using GenAI for 3.12 tasks on average, compared to 2.42 tasks for weekly user (an increase of 29%). Reported usage frequency was moderately correlated with the breadth of GenAI use across tasks (Spearman’s $\rho$ = 0.348). Compared to non-daily users, daily GenAI users were also substantially more likely to state that GenAI significantly increased their productivity, with only one respondent who used GenAI tools non-daily reported such substantial gains. Taken together, these results suggest that sporadic GenAI use is insufficient to realize meaning productivity gains and that users who perceive higher productivity gains are, in turn, more likely to use such tools often.
@@
Finding 3: The vast majority of developers (83.1%) reported using ChatGPT for SE tasks, which was nearly 2× higher than the next most popular tool (GitHub Copilot with 43.7%), suggesting that it has established itself as the leading platform for AI-assisted software development.
When interacting with GenAI tools for SE tasks, an equivalent number of participants reported using tools with a web interface (46, 63.9%) as an IDE-integrated solution (45, 62.5%). Other GenAI integrations such as GenAI APIs (14, 19.4%) and command line tools showed much lower adoption (13, 18.3%). Despite having used GenAI for development tasks, five respondents (6.9%) indicated they had not been incorporated it into their workflows.
A majority of developers (39, 54.2%) reported integrating GenAI into their workflows in two or more ways, while 28 (38.9%) reported a single integration, most commonly IDE-based (13) or web-based (12). Among participants using two integration types, 66.7% (20) reported using both web-based and IDE-integration. These findings suggest that while command-line and API-based GenAI integrations are available, they are rarely used in isolation by developers. Table 2 summarizes the # of participants across GenAI tool integrations. (No ”Other” integrations were reported by participants.)
@@
Respondents that reported daily GenAI usage also demonstrated a higher number of workflow integration types when compared with weekly users. 60.9% (31/51) reported used two or more integration methods, compared to 25% (3/12) for weekly users.
Finding 4: Developers reported no preference between incorporating web-based or IDE-integrated GenAI tools into their development workflows, but did overwhelmingly employ one or both over command-line and API-based tools. Developers who reported using GenAI tools multiple times a day were also more likely to incorporate GenAI tools into their workflows in multiple ways (e.g., Web-based, IDE-integration, etc.). This suggests that frequent GenAI users may develop more tailored or complex workflows.

### v2

# Prompting in Practice: Investigating Software Practitioners’ Use of Generative AI Tools

@@ Abstract.
The use of generative AI (GenAI) tools has fundamentally transformed software development. Central to this shift is prompt engineering, the practice of crafting textual prompts to guide GenAI tools in generating useful content. Although prompt engineering has emerged as a critical skill, prior research has focused primarily on cataloging of prompting techniques, with limited attention to how software practitioners employ GenAI within real-world development workflows. To address this gap, this study presents a systematic investigation of practitioners’ integration of GenAI tools into software development, drawing on a rigorous survey that examines prompting strategies, conversation patterns, and reliability assessments across core software development tasks.
We surveyed 72 software practitioners who actively use GenAI to characterize AI usage patterns throughout the development process. By combining qualitative and quantitative analyses of the survey responses, we identified ???? key findings that describe how prompting is performed in practice. Our study shows that while code generation is nearly universal, proficiency strongly correlates with the use of GenAI for more nuanced tasks such as debugging and code review. Practitioners also tend to favor iterative multi-turn conversations to single-shot prompting. Documentation tasks are perceived as most reliable, while complex code generation and debugging remain major challenges. Our findings provide an empirical view of practitioner practices, ranging from basic code generation to deeper integration of GenAI into development workflows, enabling us to offer recommendations for improving both GenAI tools and the ways practitioners interact with them.
@@ 1. Introduction
Generative Artificial Intelligence (GenAI) is rapidly transforming the way programmers and developers create and evolve software. With the help of GenAI tools such as GitHub Copilot (GitHub, [n. d.]), Windsurf (Cognition AI, [n. d.]), Claude Code (Anthropic, [n. d.]), and ChatGPT (OpenAI, [n. d.]), software practitioners can now easily and quickly generate new code, review and modify existing codebases, identify and correct defects, and produce code documentation. Central to these capabilities is prompting: the practice of crafting natural language instructions to guide GenAI systems toward producing useful and accurate output. Prompting has emerged as a novel engineering practice, fundamentally different from traditional software development due to its reliance on large language models (LLMs), which are inherently probabilistic and non-deterministic, but highly capable (Atil et al., 2025).
@@
Previous research has primarily focused on cataloging prompting techniques (White et al., 2023; Schulhoff et al., 2025; Sahoo et al., 2025), which offers limited information on how practitioners actually use prompting in practice: how they craft prompts for different types of software engineering (SE) tasks, how they adapt to unreliable or suboptimal output, and how GenAI is reshaping standard development activities. Previous studies have examined narrow use cases (e.g., refactoring (AlOmar et al., 2024)), specific populations (e.g., students (Alpizar-Chacon and Keuning, 2025; Choudhuri et al., 2025a)) and broader socio-organizational adoption factors (Li et al., 2024; Russo, 2024). However, there remains a lack of systematic investigation into real-world prompting practices for a variety of SE tasks, which is essential to evaluate current practices, inform practitioners and educational guidance, and drive the design of more effective GenAI technology.
@@
To address this gap, we conducted a survey of 72 software practitioners, exploring how they use GenAI tools and engage in prompting across six core SE tasks: code generation, documentation, debugging, testing, refactoring, and review. (We use software practitioners to denote professionals with experience in programming and software development, including developers, architects, technical leads, researchers, and other roles.) Our study investigates multiple dimensions of GenAI usage, including the prompting and conversational strategies that practitioners employ, how they assess and respond to the quality of GenAI output, and how prompting practices are integrated into their daily workflows. Together, these dimensions provide a holistic view of how prompting functions in modern software development and extend the current understanding of prompt engineering in practice.
@@
Our study identified ???? key findings of prompt engineering for software development. In particular, we find that while nearly all practitioners who use GenAI apply it to code generation, higher perceived proficiency strongly correlates with engagement in more types of tasks, such as code refactoring and testing. Practitioners rely primarily on iterative conversations, with daily usage critical to achieving significant productivity improvements. Our reliability assessment reveals that practitioners trust GenAI more for documentation and least for complex code generation/modification tasks. These findings have important implications for practitioners seeking to improve productivity through effective prompting strategies, for tool builders designing more reliable and intuitive GenAI, and for educators developing curricula around the evolving skill set needed for effective GenAI collaboration.
@@
This paper makes the following contributions: (1) a comprehensive characterization of how GenAI tools are used in various SE tasks; (2) empirical insights into practitioners’ prompting strategies, reliability perceptions, and common failure types; and (3) actionable implications for improving human-AI interaction in SE.
@@ 2. Study Methodology
The goal of our study is to investigate how software practitioners use prompt engineering and GenAI tools for development tasks and what challenges they face during this use. The context consists of software practitioners with experience using GenAI tools for SE.
@@
    How do software practitioners integrate GenAI tools into their development workflows?
@@
    What prompting techniques and conversation strategies do practitioners use when interacting with GenAI tools?
@@
    How reliable do practitioners perceive GenAI to be for various SE tasks and what are common issues they encounter?
Next, we detail our methodology (Figure 1), which aligns with our previous studies (Stalnaker et al., 2024a, b; Wintersgill et al., 2024; Zappin et al., 2025) and was approved by our institution’s ethics review board, including questionnaire, participant identification procedure, and data collection/distribution.
@@ 2.1. Survey Design
We designed our survey questionnaire considering the existing literature on prompt engineering and the use of GenAI tools (White et al., 2023; Schulhoff et al., 2025), following general guidelines (Groves et al., 2009) and SE-specific guidelines for survey design (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003; Linåker et al., 2015; Molléri et al., 2020). Through the survey, our goal was to build an understanding of GenAI integration across six core SE tasks: code generation, refactoring, testing, debugging, documentation, and code review. We selected these tasks because they span diverse development activities, are commonly supported by GenAI tools, and vary in complexity and risk, allowing a broad examination of prompting practices. This selection also balances task coverage with survey length, as we included task-specific questions. We implemented branching logic in the survey, so that participants were only shown questions relevant to their experiences (e.g., SE tasks for which they have used GenAI). To ensure data quality, we also included an initial question to screen out respondents with no experience using prompt-based GenAI tools. We carefully formulated the survey questions, ensuring that they were written clearly to avoid confusing or biasing the respondents. We also conducted a pilot study with graduate students from our research lab who regularly use GenAI tools. Based on their feedback, we further refined our questionnaire, improving questions that were unclear or confusing.
The survey was organized into five sections. Sec. 1 collected demographic information. Sec. 2 asked questions related to the GenAI experience of the participants, the GenAI tools they regularly use, and their GenAI development workflows. Sec. 3 and 4 investigated the use of prompting techniques and conversation strategies. The final section explored the perceived reliability of GenAI tools and the issues that practitioners face. The questions were presented in a variety of formats, including multiple-selection, Likert scales, and open text fields. The full text and logic of our survey can be found in our replication package (Otten et al., 2026).
Figure 1. Study Methodology (image credits at (Otten et al., 2026))
@@ 2.2. Participant Recruitment
Our recruitment strategy targeted software professionals with diverse backgrounds, including developers, architects, technical leads, researchers, and other roles with programming and software development experience. For survey distribution, we mainly relied on our professional networks and snowball sampling (e.g., participants sharing the survey within their networks) by directly contacting potential participants or posting participation calls on social media platforms (e.g., LinkedIn and X). To further broaden our sample and include professionals in earlier career stages, we distributed the survey to relevant organizations at our university, including the ACM chapter. Participants were not compensated monetarily or otherwise.
@@ 2.3. Data Collection and Analysis
Survey responses were collected using Qualtrics (qua, nd) between 27 May and 7 July 2025. To analyze the data, we used a mixed method approach, combining quantitative and qualitative techniques.
For multiple choice and Likert scale questions, we calculated data distributions and descriptive statistics to identify trends and patterns. Where appropriate, to facilitate comparison, we assigned ordinal Likert scale responses (e.g., ”Often” to ”Never”) to integer values (e.g., 1 to 4). This allowed us to compute mean scores to represent concepts such as the popularity of a prompting strategy or the frequency of an issue using GenAI tools.
@@
The responses to open-ended responses were analyzed using a qualitative coding approach (Spencer, 2009). Three annotators (i.e., authors) independently analyzed such responses and performed open-coding by assigning one or more codes to each response using a shared codebook (i.e., a spreadsheet). After the initial coding phase, the annotators met to collaboratively review, discuss, and reconcile the coding decisions. Through this process, the annotators resolved any discrepancies, which included differing interpretations of potentially ambiguous responses, applying the same codes in different contexts, and using broader codes rather than more specific ones. Ultimately, this process resulted in a final set of codes, definitions, and labels for each response, which are found in our replication package (Otten et al., 2026). Since codes were developed iteratively through an inductive process without an initial codebook and multiple codes could be assigned to each response, we did not base our analysis on inter-annotator agreements. We followed best open-coding practices (Spencer, 2009) and used annotator discussions to ensure the reliability of the results.
@@ 2.4. Participant Demographics
The 72 participants represent a diverse, global sample of software professionals, with reported development experience of 1-40 years (mean = 10.2, median = 10). They held a variety of roles, including researcher (64%), programmer (60%), technical lead (28%), and software architect (28%). Respondents could select multiple roles, and overlaps were common: for example, 31 of 46 participants identified as researchers (the most frequently selected role) also reported other roles, notably programmer, architect, technical lead, and/or tester. For comparative analysis, we grouped professional experience into four categories: 24 juniors (1–5 years), 22 intermediates (6–10 years), 19 seniors (11–20 years), and 7 veterans (21+ years).
@@ 3. Results
This section presents the results for each RQ, while section 4 presents a discussion and implications of these results.
@@ 3.1. RQ1: GenAI Usage Patterns
We first present the results practitioners’ GenAI usage patterns.
@@ 3.1.1. GenAI Use for SE Tasks and Proficiency
Table 1 shows that the vast majority of respondents (91.7%) regularly use GenAI for code generation. Although this might be expected, we found that debugging (47.2%) and code documentation (44.4%) emerged as the next most common tasks. The remaining tasks, testing (38.9%, refactoring (31.9%), and code review (31.9%) were less frequently reported. In general, 89% of the respondents reported using GenAI for four or fewer tasks, with two (30.6%) or three (23.6%) tasks being most common, suggesting a relatively focused use of GenAI.
When analyzing the 5-Likert point responses of self-reported proficiency in using GenAI, we found that the majority of participants (57 of 72) considered themselves proficient or very/maximally proficient (see Table 1). (Since a single veteran practitioner rated themselves as having “minimum proficiency”, we exclude this data point from future comparisons across proficiency levels.)
We also observed that proficiency correlates with the number of SE tasks for which participants use GenAI (Spearman’s $\rho$ = 0.325). Among those who reported using GenAI for four or more SE tasks, most (18 of 22, 82%) identified as very or maximally proficient. By contrast, single-task users reported being somewhat proficient or proficient (7 of 11, 64%). These results suggest that task breadth and GenAI proficiency may mutually reinforce one another.
When examining proficiency by software development experience, we found that junior and intermediate practitioners considered themselves more proficient in using GenAI that seniors and veterans. Specifically, 22 of 24 juniors and 19 of 22 intermediates reported being proficient or very/maximally proficient, compared to only 12 of 19 seniors and 4 of 7 veterans.
Finding 1: A moderate correlation between perceived GenAI proficiency and GenAI use across six SE tasks suggests that that greater proficiency is associated with broader usage, and vice versa. Junior and intermediate practitioners tend to rate themselves as more proficient in using GenAI compared to senior and veteran practitioners.
@@ 3.1.2. GenAI Usage Frequency and Tool Adoption
Respondents who reported using GenAI more frequently tended to also report using GenAI for more SE task and experiencing productivity gains (Figure 2). Daily users reported using GenAI for 3.12 tasks on average, compared to 2.42 tasks for weekly user (an increase of 29%). Reported usage frequency was moderately correlated with the breadth of GenAI use across tasks (Spearman’s $\rho$ = 0.348). Compared to non-daily users, daily GenAI users were also substantially more likely to state that GenAI significantly increased their productivity, with only one respondent who used GenAI tools non-daily reported such substantial gains. Taken together, these results suggest that sporadic GenAI use is insufficient to realize meaningful productivity gains and that users who perceive higher productivity gains are, in turn, more likely to use such tools often.
@@
Finding 3: The vast majority of practitioners (83.1%) reported using ChatGPT for SE tasks, which was nearly 2× higher than the next most popular tool (GitHub Copilot with 43.7%), suggesting that it has established itself as the leading platform for AI-assisted software development.
When interacting with GenAI tools for SE tasks, an equivalent number of participants reported using tools with a web interface (46, 63.9%) and IDE-integrated solutions (45, 62.5%). Other GenAI integrations such as GenAI APIs (14, 19.4%) and command line tools showed lower adoption (13, 18.3%). Despite having used GenAI for development tasks, five respondents (6.9%) indicated they had not been incorporated it into their workflows.
A majority of practitioners (39, 54.2%) reported integrating GenAI into their workflows in two or more ways, while 28 (38.9%) reported a single integration, most commonly IDE-based (13) or web-based (12). Among participants using two integration types, 66.7% (20) reported using both web-based and IDE-integration. These findings suggest that while command-line and API-based GenAI integrations are available, they are rarely used in isolation by practitioners. Table 2 summarizes the # of participants across GenAI tool integrations. (No ”Other” integrations were reported by participants.)
@@
Respondents that reported daily GenAI usage also demonstrated a higher number of workflow integration types when compared with weekly users: 60.9% (31/51) reported used two or more integration methods, compared to 25% (3/12) for weekly users.
Finding 4: Practitioners reported no preference between incorporating web- or IDE-based GenAI tools into their development workflows, but did overwhelmingly employ one or both over command-line and API-based tools. Those who reported using GenAI tools multiple times a day were also more likely to incorporate GenAI tools into their workflows in multiple ways (Web-based, IDE-integration, etc.). This suggests that frequent GenAI users may develop more tailored or complex workflows.